OpenStudy (anonymous):

which is equal to (3ab)0 A. 0 b. 1 c. 3 d. 3a

OpenStudy (anonymous):

@Anaise

Nnesha (nnesha):

\[\huge\rm (ab)^m = a^m *b^m\] all numbers/variables in the parentheses are raised by outside exponent and `anything to the zero power is just equal to one` \[\large\rm (anything)^0 = 1 ~~~~~(9xyz)^0 = 1 \] \[\rm (AaliyahThacker3)^0 = 1\] :P